OpenStudy (anonymous):

what is the length of the arc on a circle with radius 16ft intecepted by a 60 degree angel? round to the hundredth place pi=3.14

OpenStudy (michele_laino):

we can write this proportion: \[\Large 360:\left( {2 \cdot 3.14 \cdot 16} \right) = 60:l\] where \(l\) is the requested length

OpenStudy (anonymous):

ok im following so far

OpenStudy (michele_laino):

so, if we apply the fundamental property of proportions, we get: \[\Large l = \frac{{2 \cdot 3.14 \cdot 16}}{{360}} \cdot 60 = ...?\]

OpenStudy (anonymous):

224

OpenStudy (michele_laino):

hint: \[\Large l = \frac{{2 \cdot 3.14 \cdot 16}}{{360}} \cdot 60 = \frac{{100.48 \cdot 60}}{{360}} = ...?\]

OpenStudy (anonymous):

hold on please

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

16.746

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

oops.. the rounded value, is: \[\Large l = 16.75\;ft\]